40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3948     Entry time: Thu Nov 18 16:32:21 2010
Author: yuta 
Type: Summary 
Category: CDS 
Subject: current damping status for all optics c1sus handles 

Summary:
   I set Q-values for each ringdown of PRM, BS, ITMX, ITMY, MC1, MC2, MC3 to ~5 using QAdjuster.py.
   Here are the results;
c1susdampings.png

  Red ringdowns indicate the second try after gain setting.

Note:

  - ITMX and ITMY are referred according to MEDM screens in this entry.
  - ITMX(south) OSEM positions are currently so bad(LL and SD are all the way in/out).
        I have to change IFO_ALIGN slider values to check the damping servo. For SIDE, I couldn't do that. I reverted the slider change after the damping checking.
  - ITMY(west) somehow has opposite coil gain sign.
       Usually for the other optics, UL,UR,LR,LL is 1,-1,1,-1. But for ITMY to damp, they are -1,1,-1,1.
  - PRM damps, but ringdown doesn't look nice. There must be something funny going on.
  - SRM doesn't have OSEMs put in now.

ELOG V3.1.3-